# Maximal non-compactness of embeddings between sequence spaces

### v1

@@ Introduction
An important particular example of an operator whose compactness or non-compactness is studied is the identity. In such cases, we talk about embeddings of the corresponding spaces or classes. Measure of noncompactness of embeddings has been studied extensively, for example in connection of Sobolev embeddings and their applications to studying properties of solution to elliptic partial differential equation. The fact that here is an immense interest in studying these problems can be illustrated by the vast amount of papers including some very recent ones. To name just very few, see e.g. [3, 9, 14, 12, 13].
@@
We then turn our attention to Lorentz sequence spaces, which are collections of sequences determined by the two-parameter scale of Lorentz functionals. It is known that the Lorentz functional is always a quasinorm, which in very special cases turns to a norm or is equivalent to a norm. Lorentz spaces constitute a very interesting collection of sequence spaces, even slightly richer than their companions containing functions (note that, for example, spaces $\ell^{\infty,q}$ with finite $q$ are nontrivial).
In order to enable application of our general results to Lorentz spaces, we begin by showing that, for an arbitrary combination of $p,q\in[1,\infty]$ such that $\min\{p,q\}<\infty$, the space $\ell^{p,q}$ is a subset of $c_{0}$ by some elementary means. Next, we concentrate on characterizing the inclusions $\ell^{p_{1},q_{1}}\hookrightarrow\ell^{p_{2},q_{2}}$ between the Lorentz sequence spaces depending on the values of $p_{1},p_{2}$ and $q_{1},q_{2}$. Here, our point of departure are inequalities and inclusions mentioned in [10], which we extend and fill in details to complete the proofs and give a comprehensive picture. In most cases, we furthermore determine the exact values of the norms of the identity operators between these spaces. This knowledge is very useful in our last set of results, in which we determine whether embedding relations between Lorentz sequence spaces are maximally non-compact.
@@ 1. Preliminaries
Let $X$, $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded mapping defined on $X$ and taking values in $Y$. The ball measure of non-compactness $\alpha(T)$ of $T$ is defined as the infimum of radii $r>0$ for which there exists a finite set of balls in $Y$ of radii $r$ that covers $T(B_{X})$, where $B_{X}$ is the closed unit ball on $X$. In other words,
@@
Let us recall that given $X$ and $Y$ (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and $T\colon X\rightarrow Y$ a bounded mapping defined on $X$ and taking values in $Y$ we define the operator norm $T||$ of $T$ as follows:
@@
We will now formulate an elementary, but important inequality estimating the Ball measure of non-compactness.
@@
Let $X$ and $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded linear mapping defined on $X$ and taking values in $Y$. Then:
@@
Proof.
We will prove each of the inequalities separately:
We can see that $\alpha(T)\geq 0$ easily from the definition of the ball measure of non-compactness, because infimum of positive radii $r$ is always greater or equal to zero.
Take $\varrho>\|T\|$ arbitrarily. Then for every $x\in B_{X}$ one has $\|Tx\|\leq\|T\|<\varrho$, which in turn implies that $Tx\in B_{Y}(0,\varrho)$. Since $x$ was arbitrary, this implies that
$T(B_{X})\subset\bigcup_{i=1}^{m=1}(0+\varrho B_{Y}).$
The last inclusion is true for all $\varrho>\|T\|$. Consequently, $\alpha(T)\leq\|T\|$. ∎
@@
Let $X$ and $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded mapping defined on $X$ and taking values in $Y$. We say that $T$ is maximally non-compact if
@@
Let $X$ and $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded linear mapping defined on $X$ and taking values in $Y$. Then:
@@
    $T$ is compact if and only if for every bounded $A\subset X$ the set $T(A)$ is relatively compact in $Y$.
@@
Proof.
We will gradually prove each statement from the proposition.
  1. 1.
    $\Leftarrow$”: Let $A\subset X$ be bounded and let $\{y_{n}\}$ be a sequence in $T(A)$. Then for every $n\in\mathbb{N}$ there is $x_{n}\in A$ such that $y_{n}=T(x_{n})$. Then $\{x_{n}\}$ is a bounded sequence in $A$ and $\{y_{n}\}=\{T(y_{n})\}$ has a convergent subsequence owing to the compactness of $T$. It follows that $T(A)$ is relatively compact.
    $\Rightarrow$”: Let $A\subset X$ be bounded and and let $\{x_{n}\}$ be a sequence in $A$. Because $\{T(x_{n})\}\subset T(A)$, there exists an increasing sequence of indices $\{n_{k}\}$ such that $\{T(x_{n_{k}})\}$ converges thanks to the relative compactness of $T(A)$. So $T$ satisfies the definition of a compact operator.
  2. 2.
    Let T be compact. From (i) we have that for every bounded $A\subset X$ the set $T(A)$ is relatively compact in $Y$ therefore $T(B_{X})$ is relatively compact. Hence, $T(B_{X})$ is totally bounded. Consequently, for every $\varepsilon>0$, one has $T(B_{X})\subset\bigcup_{i=1}^{m}(y_{i}+\varepsilon B_{Y})$, in which $y_{i}\in Y$ for $i\in\{1,\dots,m\}$, and $m\in\mathbb{N}$. Therefore, $\alpha(T)=0$.
  3. 3.
    Since $\alpha(T)=0$, for every $\varepsilon>0$ we have $T(B_{X})\subset\bigcup_{i=1}^{m}(y_{i}+\varepsilon B_{Y})$, where $y_{i}\in Y$ and $m\in\mathbb{N}$, which implies that $T(B_{X})$ is totally bounded. Because $Y$ is complete, this implies that $T(B_{X})$ is relatively compact in $Y$.
    Now let $A\subset X$ be a bounded set and let $\{x_{n}\}\subset A$ be a sequence from $A$. Then there exists $r>0$ such that $\{x_{n}\}\subset A\subset B_{X}(0,r)$. Then $\frac{1}{r}x_{n}\subset B_{X}$. Now, similarly to the proof of (i), $\{T(\frac{1}{r}x_{n})\}\subset T(B_{X})$, and there exists an increasing sequence of indices $\{n_{k}\}$ such that $\{T(\frac{1}{r}x_{n_{k}})\}$ converges owing to the relative compactness of $T(B_{X})$. But then also, by the linearity of $T$ and linearity of limits, we get that $\{T(x_{n})\}=\{rT(\frac{1}{r}x_{n_{k}})\}$ converges. Therefore, for every bounded $A\subset X$ the set $T(A)$ is relatively compact in $Y$. So, from 1., $T$ is compact.
@@
Assume that $(\ell,\|\cdot\|_{\ell})$ and $(w,\|\cdot\|_{w})$ are sequence (alpha/quasi)-normed linear spaces over $\mathbb{K}$. We say that $\ell$ is a embedded into $w$ (we shall denote this as $\ell\hookrightarrow w$) if $\ell$ is a linear subspace of $w$ and the identity operator is bounded in the sense that there exists a constant $C\geq 0$ such that for every $a\in\ell$ one has $a\in w$ and $\|a\|_{w}\leq C\|a\|_{\ell}.$
@@
Let us consider the measure space $(\mathbb{N},\mathcal{P}(\mathbb{N}),m)$. For a sequence $\{a\}_{n=1}^{\infty}\subset\mathbb{K}$ we define its distribution function $m_{a}\colon[0,\infty)\rightarrow[0,\infty]$ in the following way:
@@
(alpha/quasi)-normed sequence space over $\mathbb{K}$. We say that $\ell$ is a rearrangement-invariant lattice if it satisfies the following axioms:
@@
Let $\ell$ be a (alpha/quasi)-normed sequence space over $\mathbb{K}$. We define the span $\sigma_{\ell}$ of $\ell$ as follows:
@@
We will recall the definitions of some familiar sequence spaces, namely, $\ell_{p}$ for $1\leq p<\infty$ and $c_{0}$. One has
$\ell_{p}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{K};\sum_{n=1}^{\infty}|x_{n}|^{p}<+\infty\}\text{ with norm }\|x\|_{p}=\left(\sum_{n=1}^{\infty}|x_{n}|^{p}\right)^{1/p},$
$\ell_{\infty}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{K};\sup_{n\in\mathbb{N}}|x_{n}|<+\infty\}\text{ with norm }\|x\|_{\infty}=\sup_{n\in\mathbb{N}}|x_{n}|,$
$c_{0}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{K};\;\lim_{n\to\infty}x_{n}=0\}\text{ with norm }\|x\|_{\infty}.$
@@
Let $p,q\in(0,\infty]$. We define Lorentz sequence space $\ell^{p,q}$ as a space of all sequences $a=\{a_{n}\}_{n=1}^{\infty}\subset\mathbb{K}$, for which the value $\|a\|_{p,q}$ defined below is finite:
@@
Below are some of the basic properties of Lorentz sequence spaces which follow directly from the definition:
@@
    $\|a\|_{p,q}=\left(\sum_{n=1}^{\infty}(a_{n}^{*})^{q}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}}=\|n^{\frac{1}{p}-\frac{1}{q}}a^{*}\|_{\ell^{q}}$,
@@
If $1\leq q\leq p\leq\infty$, then $\|\cdot\|_{p,q}$ is a norm. If $0<p<q\leq\infty$, $\|\cdot\|_{p,q}$ is not a norm but it is a quasi-norm satisfying, for any $a,b\in\ell^{p,q}$,
$\|a+b\|_{p,q}\leq 2^{1/p}(\|a\|_{p,q}+\|b\|_{p,q}).$
The proof can be found in [10, p. 76, Proposition 1].
@@
We will now formulate two elementary observations which will come handy in the sequel. The statements are known but, for the reader’s convenience, we recall their simple verifications.
@@
Proof.
We know that the function $f(t)=t^{q}$ is convex for $q\geq 1$, so $f(\lambda t+(1-\lambda)t)\leq\lambda f(t)+(1-\lambda)f(t)$. From this it already follows that
$\left(\frac{1}{2}\right)^{q}(a+b)^{q}=\left(\frac{1}{2}a+\frac{1}{2}b\right)^{q}\leq\frac{1}{2}(a^{q}+b^{q})$
and we are done. ∎
@@
Proof.
It is equivalent to show
$\left(\frac{a}{b}\right)^{q}+1\geq\left(\frac{a}{b}+1\right)^{q}$
We will prove this by showing that the function
$f(x)=x^{q}+1-\left(x+1\right)^{q}$
is nonnegative and increasing for $x\geq 0$. We have $f(0)=0$ and
$f^{\prime}(x)=qx^{q-1}-q\left(x+1\right)^{q-1}.$
Then $f^{\prime}(x)\geq 0$ for all $x\geq 0$, because q1<0q-1<0 and $x^{q-1}\geq\left(x+1\right)^{q-1}$. ∎
@@ 2. Main results
Let $\ell$ and $w$ be sequence (alpha/quasi)-normed linear spaces over $\mathbb{K}$ satisfying:
@@
We will now present an example to show that the Proposition 2.1 does not hold without the second condition.
@@
Let $p\in[1,\infty)$, let $\ell=\ell_{p}$ and let $w$ be a weighted $\ell_{p}$ space endowed with the norm
$\|x\|_{w}=\left(|x_{1}|^{p}+\frac{1}{2}\sum_{n=2}^{\infty}|x_{n}|^{p}\right)^{1/p}.$
Then $\ell$, $w$ and embedding $I\colon\ell\rightarrow w$ satisfy all but the second assumption of Proposition 2.1, but $I$ is not maximally non-compact and $\alpha(I)\leq\left(\frac{1}{2}\right)^{\frac{1}{p}}<\|I\|$.
@@
    One has $\ell,w\subset c_{0}$ from the necessary condition for the sum convergence.
@@
    The space $w$ does not satisfy the condition that for all $x\in w$ and for all $k\in\mathbb{N}$ it holds that $|x_{k}|\leq\|x\|_{w}$. For example take $x=e^{2}=(0,1,0,0,\dots)$, then $1=|x_{2}|>\|x\|_{w}=\left(\frac{1}{2}\right)^{\frac{1}{p}}$.
@@
    We will verify the third assumption of the Proposition 2.1.
@@
      For all $x\in\ell_{p}$ we have $\|x\|_{w}\leq\|x\|_{p}$.
    Thus, $\|I\|\leq 1$ for any $p\in[1,\infty)$. Now we will show that $I$ achieves $1$ on the closed unit ball. For $e^{1}\in B_{\ell_{p}}$ we have $\|e^{1}\|_{p}=\|e^{1}\|_{w}=1$, hence $\|I\|=1$.
We show that $\alpha(I)\leq\left(\frac{1}{2}\right)^{\frac{1}{p}}$. First we notice that for $n>1$ each member $y_{n}$ of the sequence $y\in B_{w}$ can be $2^{\frac{1}{p}}$ times larger than in $B_{\ell_{p}}$. Therefore, the ball $\left(\frac{1}{2}\right)^{\frac{1}{p}}B_{w}$ would cover all sequences $x\in B_{\ell_{p}}$ such that $x_{1}=0$. The remaining sequences $x\in B_{\ell_{p}}$, where $0<x_{1}\leq 1$ can be covered with balls $\pm e^{1}+\left(\frac{1}{2}\right)^{\frac{1}{p}}B_{w}$ and $\pm\left(\frac{1}{2}\right)^{\frac{1}{p}}e^{1}+\left(\frac{1}{2}\right)^{\frac{1}{p}}B_{w}$. Altogether, we showed that the radius in the definition of maximal non-compactness is less than or equal to $\left(\frac{1}{2}\right)^{\frac{1}{p}}$. Hence $I$ is not maximally non/compact.
@@
Let $\ell$ and $w$ be sequence (alpha/quasi)-normed linear spaces over $\mathbb{K}$ satisfying:
@@
    $\ell,\ w$ are rearrangement-invariant lattices,
  3. (3)
    the embedding operator $I\colon\ell\rightarrow w$ satisfies $0<\|I\|<\infty$, where $\|I\|=\displaystyle\sup_{x\in B_{\ell}}\|x\|_{w}$.
@@
We will prove the proposition by contradiction. Let us assume that $I$ is not maximally non-compact, then $\alpha(I)<\|I\|$. We take $\varrho\in(\alpha(I),\|I\|)$ and find $\lambda\in(0,1)$ such that:
@@
From the definition of $\|I\|$ we find $x\in B_{\ell}$ satisfying:
@@
Owing to (2.4), $\varepsilon\in c_{0}$. Then from the positive homogeneity of (alpha/quasi)-norm, our prerequisite that $w$ is a rearrangement-invariant lattice, and inequality (2.2), we get
@@
Let $\ell$ be a sequence (alpha/quasi)-normed linear space over $\mathbb{K}$ satisfying $\ell\hookrightarrow\ell^{\infty}$ and
@@
(2.11) $B_{\ell}\subset\bigcup_{i=-m}^{m}(y^{i}+\varrho B_{\ell^{\infty}}),$
@@
    Here, we get the second inequality from infy + σ_ℓ ∈(λ_k-1 + σ_ℓ/2, λ_k + σ_ℓ/2]. On the other hand, using the definition of $\lambda_{k}$ and inequality (2.10), we arrive at
@@
Altogether, we showed that $y\in y^{k}+\varrho B_{\ell^{\infty}}$, and (2.11) follows. Therefore $\alpha(I)\leq\sigma_{\ell}/2<\|I\|$ and $I$ is not maximally non-compact. ∎
@@
$B_{c_{0}}\subset\bigcup_{i=1}^{m}(y^{i}+\varrho B_{\ell_{\infty}}).$

### v2

@@ Introduction
An important particular example of an operator whose compactness or non-compactness is studied is the identity. In such cases, we talk about embeddings of the corresponding spaces or classes. Measure of noncompactness of embeddings has been studied extensively, for example in connection with Sobolev embeddings and their applications to studying properties of solution to elliptic partial differential equation. The fact that there is an immense interest in studying these problems can be illustrated by the vast amount of papers including some very recent ones. To name just very few, see e.g. [3, 9, 12, 13, 14].
@@
We then turn our attention to Lorentz sequence spaces, which are collections of sequences determined by the two-parameter scale of Lorentz functionals. It is known that the Lorentz functional is always a quasinorm, which in some special cases turns to a norm or is equivalent to a norm. Lorentz spaces constitute a very interesting collection of sequence spaces, even slightly richer than their companions containing functions (note that, for example, spaces $\ell^{\infty,q}$ with finite $q$ are nontrivial). Let us recall that Lorentz function spaces have been treated in various monographs, see e.g. [2, 15, 18].
In order to enable application of our general results to Lorentz spaces, we begin by showing that, for an arbitrary combination of $p,q\in[1,\infty]$ such that $\min\{p,q\}<\infty$, the space $\ell^{p,q}$ is a subset of $c_{0}$ by some elementary means. Next, we concentrate on characterizing the inclusions $\ell^{p_{1},q_{1}}\hookrightarrow\ell^{p_{2},q_{2}}$ between the Lorentz sequence spaces depending on the values of $p_{1},p_{2}$ and $q_{1},q_{2}$. Here, our point of departure are inequalities and inclusions mentioned in [10], which we extend and fill in detail to complete the proofs and give a comprehensive picture. In most cases, we furthermore determine the exact values of the norms of the identity operators between these spaces. This knowledge is very useful in our last set of results, in which we determine whether embedding relations between Lorentz sequence spaces are maximally non-compact.
@@ 1. Preliminaries
Let $X$, $Y$ be (quasi)-normed linear spaces over $\mathbb{R}$ and let $T\colon X\rightarrow Y$ be a bounded mapping defined on $X$ and taking values in $Y$. The ball measure of non-compactness $\alpha(T)$ of $T$ is defined as the infimum of radii $r>0$ for which there exists a finite set of balls in $Y$ of radii $r$ that covers $T(B_{X})$, where $B_{X}$ is the closed unit ball on $X$. In other words,
@@
Let us recall that given $X$ and $Y$ (quasi)-normed linear spaces over $\mathbb{R}$ and $T\colon X\rightarrow Y$ a bounded mapping defined on $X$ and taking values in $Y$ we define the operator norm $T||$ of $T$ as follows:
@@
We will now formulate an elementary, but important inequality estimating the ball measure of non-compactness
@@
Let $X$ and $Y$ be (quasi)-normed linear spaces over $\mathbb{R}$ and let $T\colon X\rightarrow Y$ be a bounded linear mapping defined on $X$ and taking values in $Y$. Then:
@@
The proof directly follows from the definition of the ball measure of non-compactness.
@@
Let $X$ and $Y$ be (quasi)-normed linear spaces over $\mathbb{R}$ and let $T\colon X\rightarrow Y$ be a bounded mapping defined on $X$ and taking values in $Y$. We say that $T$ is maximally non-compact if
@@
Let $X$ and $Y$ be (quasi)-normed linear spaces over $\mathbb{R}$ and let $T\colon X\rightarrow Y$ be a bounded linear mapping defined on $X$ and taking values in $Y$. Then:
@@
    If $T$ is compact, then $\alpha(T)=0$.
@@
Variants of Proposition 1.4 are well known. The assertion can be easily obtained using standard results in analysis, in particular, the definition of a compact linear operator and the equivalence of total boundedness and relative compactness in complete metric spaces.
@@
Assume that $(\ell,\|\cdot\|_{\ell})$ and $(w,\|\cdot\|_{w})$ are sequence (quasi)-normed linear spaces over $\mathbb{R}$. We say that $\ell$ is embedded into $w$ (we shall denote this as $\ell\hookrightarrow w$) if $\ell$ is a linear subspace of $w$ and the identity operator is bounded in the sense that there exists a constant $C\geq 0$ such that for every $a\in\ell$ one has $a\in w$ and $\|a\|_{w}\leq C\|a\|_{\ell}.$
@@
Let us consider the measure space $(\mathbb{N},\mathcal{P}(\mathbb{N}),m)$. For a sequence $\{a\}_{n=1}^{\infty}\subset\mathbb{R}$ we define its distribution function $m_{a}\colon[0,\infty)\rightarrow[0,\infty]$ in the following way:
@@
(quasi)-normed sequence space over $\mathbb{R}$. We say that $\ell$ is a rearrangement-invariant lattice if it satisfies the following axioms:
@@
Let $\ell$ be a (quasi)-normed sequence space over $\mathbb{R}$. We define the span $\sigma_{\ell}$ of $\ell$ as follows:
@@
We will recall the definitions of some familiar sequence spaces, namely, $\ell^{p}$ for $1\leq p\leq\infty$ and $c_{0}$. One has
[…]
$\ell_{\infty}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{R};\sup_{n\in\mathbb{N}}|x_{n}|<+\infty\}\text{ with norm }\|x\|_{\infty}=\sup_{n\in\mathbb{N}}|x_{n}|\text{ and }p=\infty,$
$c_{0}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{R};\;\lim_{n\to\infty}x_{n}=0\}\text{ with norm }\|x\|_{\infty}.$
@@
Let $p,q\in(0,\infty]$. We define Lorentz sequence space $\ell^{p,q}$ as a space of all sequences $a=\{a_{n}\}_{n=1}^{\infty}\subset\mathbb{R}$, for which the value $\|a\|_{p,q}$ defined below is finite:
@@
Below are some of the basic properties of Lorentz sequence spaces which follow directly from the definition for all $p,q\in(0,\infty]$:
@@
    $\|a\|_{p,q}=\|n^{\frac{1}{p}-\frac{1}{q}}a_{n}^{*}\|_{\ell^{q}}$,
@@
Standard techniques show that $\|\cdot\|_{p,q}$ is a quasinorm whenever $p,q\in(0,\infty]$. Moreover, it can be shown that, for $1\leq q\leq p\leq\infty$, it is a norm. For further details, see e.g. [10, p. 76, Proposition 1].
@@
We will now formulate two well-known elementary observations which will come handy in the sequel.
@@ 2. Main results
Let $\ell$ and $w$ be sequence (quasi)-normed linear spaces over $\mathbb{R}$ satisfying:
@@
We will now present an example which demonstrates that Proposition 2.1 does not hold without the second condition.
@@
Let $p\in[1,\infty)$ and $\alpha\in(0,1)$. Set $\ell=\ell^{p}$ and let $w_{\alpha}$ be the weighted $\ell^{p}$ space endowed with the norm
$\|x\|_{w_{\alpha}}=\left(|x_{1}|^{p}+\alpha\sum_{n=2}^{\infty}|x_{n}|^{p}\right)^{1/p}.$
Then $\ell$, $w_{\alpha}$ and embedding $I\colon\ell\rightarrow w_{\alpha}$ satisfy all but the second assumption of Proposition 2.1, but $I$ is not maximally non-compact and $\alpha(I)\leq\alpha^{\frac{1}{p}}<\|I\|$.
@@
    One has $\ell,w_{\alpha}\subset c_{0}$ owing to the necessary condition for the sum convergence.
@@
    The space $w_{\alpha}$ does not satisfy the condition that for all $x\in w_{\alpha}$ and for all $k\in\mathbb{N}$ it holds that $|x_{k}|\leq\|x\|_{w_{\alpha}}$. For example take $x=e^{2}=(0,1,0,0,\dots)$, then $1=|x_{2}|>\|x\|_{w_{\alpha}}=\alpha^{\frac{1}{p}}$.
@@
    We will verify the third assumption of Proposition 2.1.
@@
      For all $x\in\ell^{p}$ we have $\|x\|_{w_{\alpha}}\leq\|x\|_{p}$ and $\|e^{1}\|_{p}=\|e^{1}\|_{w_{\alpha}}=1$, hence $\|I\|=1$.
We show that $\alpha(I)\leq\alpha^{\frac{1}{p}}$. First we notice that for $n>1$ each element $y_{n}$ of a sequence $y\in B_{w_{\alpha}}$ can be taken $\left(\frac{1}{\alpha}\right)^{\frac{1}{p}}$ times larger than in $B_{\ell^{p}}$. Therefore, the ball $\alpha^{\frac{1}{p}}B_{w_{\alpha}}$ would cover all sequences $x\in B_{\ell^{p}}$ such that $x_{1}=0$. The remaining sequences $x\in B_{\ell^{p}}$, where $0<x_{1}\leq 1$, can be covered by a finite number of balls of radius $\alpha^{\frac{1}{p}}$, where the number of balls depends on the size of $\alpha^{\frac{1}{p}}$. More precisely, if $\alpha^{\frac{1}{p}}\in\left(\frac{1}{k+1},\frac{1}{k}\right]$ for some $k\in\mathbb{N}$, we can cover the remaining sequences with $(2k+1)$ balls of type
$\left(0+\alpha^{\frac{1}{p}}B_{w_{\alpha}}\right),\left(\pm\alpha^{\frac{1}{p}}e^{1}+\alpha^{\frac{1}{p}}B_{w_{\alpha}}\right),\left(\pm 2\alpha^{\frac{1}{p}}e^{1}+\alpha^{\frac{1}{p}}B_{w_{\alpha}}\right),\dots,\left(\pm(k-1)\alpha^{\frac{1}{p}}e^{1}+\alpha^{\frac{1}{p}}B_{w_{\alpha}}\right).$
Altogether, we we have shown that the radius in the definition of maximal non-compactness is less than or equal to $\alpha^{\frac{1}{p}}$. Hence $I$ is not maximally non-compact.
@@
Let $\ell$ and $w$ be sequence (quasi)-normed linear spaces over $\mathbb{R}$ satisfying:
@@
    $\ell,\ w$ are rearrangement-invariant lattices.
@@
Let us assume $\|I\|>0$, since the theorem holds for the pathological case when $\ell$ is a trivial space. We will prove the proposition by contradiction. Let us assume that $I$ is not maximally non-compact, then $\alpha(I)<\|I\|$. We take $\varrho\in(\alpha(I),\|I\|)$ and find $\lambda\in(0,1)$ such that:
@@
From the definition of $\|I\|=\displaystyle\sup_{x\in B_{\ell}}\|x\|_{w}$ we find $x\in B_{\ell}$ satisfying:
@@
Owing to (2.4), $\varepsilon\in c_{0}$. Then from the positive homogeneity of a (quasi)-norm, our prerequisite that $w$ is a rearrangement-invariant lattice, and inequality (2.2), we get
@@
Let $\ell$ be a sequence (quasi)-normed linear space over $\mathbb{R}$ satisfying $\ell\hookrightarrow\ell^{\infty}$ and
@@
(2.11) $I(B_{\ell})\subset\bigcup_{i=-m}^{m}(y^{i}+\varrho B_{\ell^{\infty}}),$
@@
    Here, we get the second inequality from
    $\inf y+\sigma_{\ell}\in(\lambda_{k-1}+\sigma_{\ell}/2,\lambda_{k}+\sigma_{\ell}/2].$
    On the other hand, using the definition of $\lambda_{k}$ and inequality (2.10), we arrive at
@@
Altogether, we showed that $y\in y^{k}+\varrho B_{\ell^{\infty}}$ for some $k\in\{-m,\dots,m\}$, and (2.11) follows. Therefore $\alpha(I)\leq\sigma_{\ell}/2<\|I\|$ and $I$ is not maximally non-compact. ∎
@@
$I(B_{c_{0}})\subset\bigcup_{i=1}^{m}(y^{i}+\varrho B_{\ell_{\infty}}).$